# Role of octupole shape degree of freedom in neutron-rich odd-mass xenon isotopes

### v1

@@ Abstract
Influences of the octupole shape degree of freedom on low-energy spectra of neutron-rich odd-mass xenon isotopes are studied within the interacting boson-fermion model that is based on the nuclear density functional theory. The interacting-boson Hamiltonian describing low-energy quadrupole and octupole collective states of the even-even nuclei 140,142,144Xe, single-particle energies and occupation probabilities for an unpaired neutron in the odd-mass nuclei 141,143,145Xe, are determined based on the axially-symmetric quadrupole-octupole deformation-constrained self-consistent mean-field calculations with a choice of the energy density functional and pairing interaction. Strength parameters of the boson-fermion interactions are empirically determined to reproduce a few low-lying levels of each odd-mass nucleus. The mean-field calculation predicts for 142Xe a potential energy surface that is notably soft in the octupole deformation with a non-zero octupole global minimum. The octupole correlations are shown to be relevant in positive-parity excited states of 143,145Xe.
@@ I Introduction
Search for the reflection asymmetric, octupole deformation in nuclei has been a recurrent theme of interest that is under active investigations for several decades. The octupole correlations are considered to be enhanced in several specific mass regions corresponding to the neutron $N$ and/or proton $Z$ numbers equal to 34, 56, 88, 134, $\ldots$, in which coupling between the normal and unique-parity single-particle orbitals differing in the total and orbital angular momenta by $3\hbar$ is possible [1]. Most of those nuclei with the aforementioned nucleon numbers are short-lived, but have recently become accessible by experiments using radioactive beams. In addition, since observation of a large electric dipole moment in the octupole-deformed nucleus would indicate violation of time-reversal (T) or charge-parity (CP) symmetry, it is also relevant to a possible extension of the standard model of particle physics [2]. Evidence for a static octupole deformation was found in light actinide region, i.e., 220Rn and 224Ra, by the experiment at ISOLDE, CERN [3]. The experiments carried out at Argonne National Laboratory found another candidates for the octupole deformation in the neutron-rich region, i.e., 144Ba [4] and 146Ba [5].
@@
The theoretical method employed in the present study is the interacting boson model (IBM) that is based on the nuclear density functional theory. The starting point is the constrained self-consistent mean-field (SCMF) calculations with a choice of the energy density functional (EDF) and pairing interaction, providing the potential energy surface (PES) as a function of the axially-symmetric quadrupole and octupole deformations. The quadrupole-octupole SCMF PES is then mapped onto the expectation value of the Hamiltonian of the IBM, that consists of monopole $s$ (with spin and parity $J^{\pi}=0^{+}$), quadrupole $d$ ($J^{\pi}=2^{+}$), and octupole $f$ ($J^{\pi}=3^{-}$) bosons. The mapping procedure specifies the $sdf$-IBM Hamiltonian describing low-lying positive- and negative-parity states of even-even nuclei. Odd-mass nuclei are dealt with by means of the particle-boson coupling in the interacting bosom-fermion model (IBFM) [9], with the even-even $sdf$-IBM core Hamiltonian determined by the mapping procedure. The same SCMF calculations provide microscopic input to determine most parts of the $sdf$-IBFM Hamiltonian. Strength parameters for the boson-fermion interactions are, however, obtained empirically to reproduce low-energy spectra of each odd-mass nucleus.
@@ II Theoretical framework
The constrained SCMF calculations are carried out within the relativistic Hartree-Bogoliubov (RHB) model [22, 23] using the density-dependent point-coupling (DD-PC1) [24] functional for particle-hole channel and a separable pairing force of finite range [25] in the particle-particle channel. The constraints are those on the axially-symmetric mass quadrupole $Q_{20}$ and octupole $Q_{30}$ moments, which are related to the dimensionless shape variables $\beta_{2}$ and $\beta_{3}$, respectively:
@@
where the first, second, and third terms on the right-hand side represent the $sdf$-IBM Hamiltonian, single-fermion Hamiltonian, and boson-fermion interaction, respectively. For $\hat{H}_{\rm B}$ we consider the form
@@
Figure 1: Axially-symmetric quadrupole ($\beta_{2}$) - octupole ($\beta_{3}$) constrained PESs for the even-even 140,142,144Xe isotopes, computed by the RHB method using the DD-PC1 functional and a separable pairing force of finite range. Total SCMF energies are plotted up to 2 MeV with respect to the global minimum, indicated by the open triangle. Energy difference between neighboring contours is 100 keV.
@@ III Results for even-even Xe
Figure 1 shows the RHB-SCMF PESs in terms of the axially-symmetric quadrupole $\beta_{2}$ and octupole $\beta_{3}$ deformations for the even-even nuclei 140,142,144Xe, plotted up to the 2 MeV from the global minimum. All these nuclei exhibit prolate deformation along the $\beta_{2}$ deformation with the equilibrium minimum found at $\beta_{2}\approx 0.15$, 0.20, and 0.20 for 140Xe, 142Xe, 144Xe, respectively. Notable features are that for 142Xe ($N=88$) the global minimum is found at a non-zero octupole deformation, $(\beta_{2},\beta_{3})=(0.2,0.05)$, and that the potential is soft along $\beta_{3}$ deformation, as the energy at the global minimum is only 32 keV lower than the lowest energy along the $\beta_{3}=0$ axis. These characteristics of the PESs indicate the relevance of the octupole correlations in the 142Xe nucleus, which also confirms that the octupole deformation is most pronounced at $N=88$. In the previous study of Ref. [16], the RHB-SCMF calculation for the 144Ba nucleus also suggested a $\beta_{3}$-soft potential but with a much steeper $\beta_{3}\neq 0$ global minimum at $(\beta_{2},\beta_{3})\approx(0.2,0.1)$. For 144Xe, the non-zero octupole minimum is no longer seen, but the potential is still soft in the $\beta_{3}$ deformation, to a similar extent to that for the 142Xe.
@@
Figure 5: Evolution of low-lying negative-parity [$\pi=-1$, panels (a) and (b)] and positive-parity [$\pi=+1$, panels (c) and (d)] spectra for the odd-mass nuclei 141,143,145Xe as functions of $N$. The calculated results from the $sdf$-IBFM are shown in panels (a) and (c), and the experimental data [18] are depicted in panels (b) and (d).
@@ IV.2 Detailed energy spectrum
The present calculation, however, gives the $\Delta I=2$ positive-parity band built on the ${13/2}^{+}_{1}$ level, with the bandhead level in agreement with experiment. The predicted ${13/2}^{+}_{1}$ band also exhibits large $E2$ transitions: $B(E2;{17/2}^{+}_{1}\to{13/2}^{+}_{1})=39$ W.u., and $B(E2;{21/2}^{+}_{1}\to{17/2}^{+}_{1})=38$ W.u. Experimentally a $\Delta I=2$ band built on the ${15/2}^{+}$ level at 1333 keV is observed. At about the same excitation energy a band that starts from the ${11/2}^{+}_{1}$ state is obtained within the $sdf$-IBFM, which is connected by the $E2$ transitions of $B(E2;{15/2}^{+}_{1}\to{11/2}^{+}_{1})=27$ W.u., and $B(E2;{19/2}^{+}_{1}\to{15/2}^{+}_{1})=30$ W.u.
@@
Figure 8 compares the $sdf$-IBFM energy spectra calculated with and without taking into account the $\beta_{3}\neq 0$ global minimum in the mapping procedure for the even-even core 142Xe. The same boson-fermion strength parameters are used for the two different $sdf$-IBFM calculations, the only difference being in the Hamiltonian parameters for the even-even core. It is suggested that the negative-parity states are not affected by the assumption of the zero octupole minimum in the mapping, since the negative-parity states are mainly comprised of normal-parity single-particle configurations coupled to the $sd$-boson space, whereas the $f$-boson contributions are minor, $\braket{\hat{n}_{f}}\approx 0.5$ [cf. Table 7]. On the other hand, the positive-parity levels calculated with the $\beta_{3}=0$ minimum are higher than those with the $\beta_{3}\neq 0$ minimum, since the the positive-parity states for 143Xe are mostly of octupole nature, i.e., $\braket{\hat{n}_{f}}\approx 1.5$, and are affected by the presence of the $\beta_{3}\neq 0$ minimum.
@@ Acknowledgements.
The author thanks Atsuko Odahara and Kenichi Yoshida for valuable discussions. This work has been supported in part by the RCNP Collaboration Research Network program as the project number COREnet-051.

### v2

@@ Abstract
Influences of the octupole shape degree of freedom on low-energy spectra of neutron-rich odd-mass xenon isotopes are studied within the interacting boson-fermion model that is based on the nuclear density functional theory. The interacting-boson Hamiltonian describing low-energy quadrupole and octupole collective states of the even-even nuclei 140,142,144Xe, single-particle energies, and occupation probabilities for an unpaired neutron in the odd-mass nuclei 141,143,145Xe, are determined based on the axially symmetric quadrupole-octupole deformation-constrained self-consistent mean-field calculations with a choice of the energy density functional and pairing interaction. Strength parameters of the boson-fermion interactions are empirically determined to reproduce a few low-lying levels of each odd-mass nucleus. The mean-field calculation predicts for 142Xe a potential energy surface that is notably soft in the octupole deformation with a non-zero octupole global minimum. The octupole correlations are shown to be relevant in positive-parity excited states of 143,145Xe.
@@ I Introduction
Search for the reflection asymmetric, octupole deformation in nuclei has been a recurrent theme of interest that is under active investigations for several decades. The octupole correlations are considered to be enhanced in several specific mass regions corresponding to the neutron $N$ and/or proton $Z$ numbers equal to 34, 56, 88, 134, $\ldots$, in which coupling between the normal and unique-parity single-particle orbitals differing in the total and orbital angular momenta by $3\hbar$ is possible [1]. Most of those nuclei with the aforementioned nucleon numbers are short lived, but have recently become accessible by experiments using radioactive beams. In addition, since observation of a large electric dipole moment in the octupole-deformed nucleus would indicate violation of time-reversal (T) or charge-parity (CP) symmetry, it is also relevant to a possible extension of the standard model of particle physics [2]. Evidence for a static octupole deformation was found in light actinide region, i.e., 220Rn and 224Ra, by the experiment at ISOLDE, CERN [3]. The experiments carried out at Argonne National Laboratory found another candidates for the octupole deformation in the neutron-rich region, i.e., 144Ba [4] and 146Ba [5].
@@
The theoretical method employed in the present study is the interacting boson model (IBM) that is based on the nuclear density functional theory. The starting point is the constrained self-consistent mean-field (SCMF) calculations with a choice of the energy density functional (EDF) and pairing interaction, providing the potential energy surface (PES) as a function of the axially symmetric quadrupole and octupole deformations. The quadrupole-octupole SCMF PES is then mapped onto the expectation value of the Hamiltonian of the IBM, that consists of monopole $s$ (with spin and parity $J^{\pi}=0^{+}$), quadrupole $d$ ($J^{\pi}=2^{+}$), and octupole $f$ ($J^{\pi}=3^{-}$) bosons. The mapping procedure specifies the $sdf$-IBM Hamiltonian describing low-lying positive- and negative-parity states of even-even nuclei. Odd-mass nuclei are dealt with by means of the particle-boson coupling in the interacting bosom-fermion model (IBFM) [9], with the even-even $sdf$-IBM core Hamiltonian determined by the mapping procedure. The same SCMF calculations provide microscopic input to determine most parts of the $sdf$-IBFM Hamiltonian. Strength parameters for the boson-fermion interactions are, however, obtained empirically to reproduce low-energy spectra of each odd-mass nucleus.
@@ II Theoretical framework
The constrained SCMF calculations are carried out within the relativistic Hartree-Bogoliubov (RHB) model [22, 23] using the density-dependent point-coupling (DD-PC1) [24] functional for particle-hole channel and a separable pairing force of finite range [25] in the particle-particle channel. The constraints are those on the axially symmetric mass quadrupole $Q_{20}$ and octupole $Q_{30}$ moments, which are related to the dimensionless shape variables $\beta_{2}$ and $\beta_{3}$, respectively:
@@
where the first, second, and third terms on the right-hand side represent the $sdf$-IBM Hamiltonian, single-fermion Hamiltonian, and boson-fermion interaction, respectively. Here $\hat{H}_{\rm B}$ takes the form
@@
Figure 1: Axially symmetric quadrupole ($\beta_{2}$) - octupole ($\beta_{3}$) constrained PESs for the even-even 140,142,144Xe isotopes, computed by the RHB method using the DD-PC1 functional and a separable pairing force of finite range. Total SCMF energies are plotted up to 2 MeV with respect to the global minimum, indicated by the open triangle. Energy difference between neighboring contours is 100 keV.
@@ III Results for even-even Xe
Figure 1 shows the RHB-SCMF PESs in terms of the axially symmetric quadrupole $\beta_{2}$ and octupole $\beta_{3}$ deformations for the even-even nuclei 140,142,144Xe, plotted up to the 2 MeV from the global minimum. All these nuclei exhibit prolate deformation along the $\beta_{2}$ deformation with the equilibrium minimum found at $\beta_{2}\approx 0.15$, 0.20, and 0.20 for 140Xe, 142Xe, 144Xe, respectively. Notable features are that for 142Xe ($N=88$) the global minimum is found at a non-zero octupole deformation, $(\beta_{2},\beta_{3})=(0.2,0.05)$, and that the potential is soft along $\beta_{3}$ deformation, as the energy at the global minimum is only 32 keV lower than the lowest energy along the $\beta_{3}=0$ axis. These characteristics of the PESs indicate the relevance of the octupole correlations in the 142Xe nucleus, which also confirms that the octupole deformation is most pronounced at $N=88$. In the previous study of Ref. [16], the RHB-SCMF calculation for the 144Ba nucleus also suggested a $\beta_{3}$-soft potential but with a much steeper $\beta_{3}\neq 0$ global minimum at $(\beta_{2},\beta_{3})\approx(0.2,0.1)$. For 144Xe, the non-zero octupole minimum is no longer seen, but the potential is still soft in the $\beta_{3}$ deformation, to a similar extent to that for the 142Xe.
@@
Figure 5: Evolution of low-lying negative-parity [$\pi=-1$, (a) and (b)] and positive-parity [$\pi=+1$, (c) and (d)] spectra for the odd-mass nuclei 141,143,145Xe as functions of $N$. The calculated results from the $sdf$-IBFM are shown in (a) and (c), and the experimental data [18] are depicted in (b) and (d).
@@ IV.2 Detailed energy spectrum
The present calculation, however, gives the $\Delta I=2$ positive-parity band built on the ${13/2}^{+}_{1}$ level, with the band-head level in agreement with experiment. The predicted ${13/2}^{+}_{1}$ band also exhibits large $E2$ transitions: $B(E2;{17/2}^{+}_{1}\to{13/2}^{+}_{1})=39$ W.u., and $B(E2;{21/2}^{+}_{1}\to{17/2}^{+}_{1})=38$ W.u. Experimentally a $\Delta I=2$ band built on the ${15/2}^{+}$ level at 1333 keV is observed. At about the same excitation energy a band that starts from the ${11/2}^{+}_{1}$ state is obtained within the $sdf$-IBFM, which is connected by the $E2$ transitions of $B(E2;{15/2}^{+}_{1}\to{11/2}^{+}_{1})=27$ W.u., and $B(E2;{19/2}^{+}_{1}\to{15/2}^{+}_{1})=30$ W.u.
@@
Figure 8 compares the $sdf$-IBFM energy spectra calculated with and without taking into account the $\beta_{3}\neq 0$ global minimum in the mapping procedure for the even-even core 142Xe. The same boson-fermion strength parameters are used for the two different $sdf$-IBFM calculations, the only difference being in the Hamiltonian parameters for the even-even core. It is suggested that the negative-parity states are not affected by the assumption of the zero octupole minimum in the mapping, since the negative-parity states are mainly comprised of normal-parity single-particle configurations coupled to the $sd$-boson space, whereas the $f$-boson contributions are minor, $\braket{\hat{n}_{f}}\approx 0.5$ [cf. Table 7]. On the other hand, the positive-parity levels calculated with the $\beta_{3}=0$ minimum are higher than those with the $\beta_{3}\neq 0$ minimum, since the positive-parity states for 143Xe are mostly of octupole nature, i.e., $\braket{\hat{n}_{f}}\approx 1.5$, and are affected by the presence of the $\beta_{3}\neq 0$ minimum.
@@ Acknowledgements.
The author thanks Atsuko Odahara and Kenichi Yoshida for valuable discussions. This work has been supported in part by the RCNP Collaboration Research Network program as the Project No. COREnet-051.